OpenStudy (anonymous):

Please help. Find f(antiderivative) . f'(t) =6 cos t + sec^2t, -pi/2pi/2, f(pi/3)=4

OpenStudy (anonymous):

\[f'(t)=6\cos t+\sec^2t~~\Rightarrow~~\int f'(t)~dt=\int(6\cos t+\sec^2t)~dt\] Integrating gives \[f(t)=6\sin t+\tan t+C\] You can find the value of \(C\) using the given initial value, \(f\left(\dfrac{\pi}{3}\right)=4\): \[4=6\sin\frac{\pi}{3}+\tan\frac{\pi}{3}+C\]